OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (3, 6) and (–2, 1)?

OpenStudy (anonymous):

find the slope m=(6-1)/(3-(-2)=5/5=1 now use point slope form y-y0=m(x-x0) y-6=1(x-3) y-6=x-3